OpenStudy (anonymous):

Use the normal approximation to the binomial to find that probability for the specific value of X. n=30, p=0.6, X=17 A) 0.14 B) 0.57 C) 0.34 D) 0.6

OpenStudy (kropot72):

Do you know the formulas for the mean and the standard deviation of the normal approximation to the binomial distribution?

OpenStudy (anonymous):

I think so. n*p is the mean, and the standard deviation is the square root of the mean? :/

OpenStudy (anonymous):

sorry, i'm just rusty because i'm making up this test and we learned the stuff about a week ago.. i appreciate the help tho!!

OpenStudy (kropot72):

\[mean=n \times p\] \[standard\ deviation=\sqrt{npq}\] where q =1 - p So can you find the values of the mean and the standard deviation and post them please?

OpenStudy (anonymous):

sure.. one second my friend

OpenStudy (anonymous):

mean=18 and standard dev.= 2.68 ??

OpenStudy (anonymous):

is that with the chart?

OpenStudy (kropot72):

Good work. Now you need to find the z-score for X = 17 to enable you to use the standard normal distribution table. Do you know how to find the z-score? You need to use the formula\[z=\frac{X-\mu}{\sigma}\]

OpenStudy (anonymous):

wait.. how am i finding the z-score?

OpenStudy (anonymous):

nvm lol

OpenStudy (anonymous):

-0,37313433

OpenStudy (kropot72):

You are reasonably correct. However I have just read the question more carefully and realise that the probability of a discrete value of X (17) is required. Please wait a few minutes.

OpenStudy (anonymous):

would it be c) 0.34??

OpenStudy (kropot72):

We are changing from a discrete distribution (binomial) to a continuous distribution (normal). We need to find the following probability:\[P(16\le X \le18)\] To do this we need to find two z-scores each of which has a correction term 0.5 caused by the change from a discrete to a continuous distribution. \[z _{\alpha}=\frac{16-18-0.5}{2.683}=you\ can\ calculate\] \[z _{\beta}=\frac{18-18+0.5}{2.683}=you\ can\ calculate\] Please post the calculated values. Then we can go to the next and final step.

OpenStudy (anonymous):

-o.93179277 and 0.18635855

OpenStudy (kropot72):

Excellent! Now we need to use a standard normal distribution table to find the two probability values corresponding to these z-scores. I suggest that you use a table found by going to this link http://www.math.bgu.ac.il/~ngur/Teaching/probability/ and selecting 'normal.pdf' Please post the two values.

OpenStudy (anonymous):

0.1762 and 0.5753

OpenStudy (kropot72):

Finally we need to find the difference between these two values of probability:\[P(X=17)=0.5753-0.1762\approx 0.4\] This value is not one of the given choices. Therefore I will calculate by using the binomial distribution and compare the result. Please wait.

OpenStudy (kropot72):

The binomial distribution calculation gives P(X = 17) = 0.136. Please wait a few minutes and I will recheck the normal approximation method.

OpenStudy (anonymous):

dont worry about it man, youve done enough. im sure ill get some partial credit and imma bout to go to bed. you think you could help me with one last question that might be easier? if its another tough one, then dont worry about it buddy

OpenStudy (anonymous):

the average diameter of sand dollars on a certain island is 4.50 centimeters with a standard deviation of 1.00 centimeters. if 9 sand dollars are chosen at random for a collection, find the probability that the average diameter of those sand dollars is more than 4.10 centimeters. assume that the variable is normally distributed.

OpenStudy (anonymous):

a. 0.615 b. 0.820 c. 0.385 d. 0.885

OpenStudy (anonymous):

just let me know what you want to do, because im getting tired lol.. i really appreciate all of your help though!!!!!

OpenStudy (kropot72):

I will look at this last one. It should be quick :)

OpenStudy (anonymous):

ok thanks a lot!!

OpenStudy (anonymous):

im getting off in about 3 mins kropot. so if its taking you a while, don't worry about it buddy!!

OpenStudy (kropot72):

Sorry I can't be sure it will be quick :( Have a good sleep!

OpenStudy (anonymous):

don't be sorry at all. you too, good night!!!!!

OpenStudy (kropot72):

Regarding the first question, the correct answer is obtained if the values of the two z-scores are calculated as follows: \[z _{\alpha}=\frac{16-18+0.5}{2.683}=-0.56\] \[z _{\beta}=\frac{18-18-0.5}{2.683}=-0.186\] If you use the table at the link that I gave previously, find the two probability values, and then find the difference: \[P(X=17)=P _{\beta}-P _{\alpha}\]